FIX = (X+4)*(x-2)} (x-3) Find the intercept) at whichf crosses The X axis Express in ordered pairs.
Answer #1

Sol: y -0. x-intere cept means -> (x+4)2(x-2)} (x-3) = 0 > x+4 =004 x 2 = oor x-3=0 x = -4 or x 2 or x=3 so the x-intercepts